OpenStudy (shesolitt):

I WILL MEDAL AND FAN What is the slope-intercept form equation of the line that passes through (1, 3) and (3, 7)? y = −2x + 1 y = −2x − 1 y = 2x + 1 y = 2x − 1

OpenStudy (shesolitt):

@helpppp @Vuriffy @Will.H

OpenStudy (superpanda1022):

did you draw it on graphing paper yet?

OpenStudy (superpanda1022):

let me help you out here

OpenStudy (will.h):

Find the slope 1st using Y2-Y1/X2-X1 Substitute 7-3/3-1 Slope = 4/2 Slope = 2 Now use this form Y - y1 = m(x - x1) Use this point (1,3) Substitute Y - 3 = 2(x - 1) Y - 3 = 2x -2 Isolate y Y = 2x +1 Hope that helps

OpenStudy (superpanda1022):

...

OpenStudy (shesolitt):

thank you @Will.H (:

OpenStudy (helpppp):

simply fill in for your x and y values, the one that is true for both is your answer

OpenStudy (will.h):

You welcome